Can you help me solve: x + ( x - 6) = 94
Answer by waynest(281)

x + (x -6) = 94
2x - 6 + 6 = 94 + 6
2x = 100
2x/2 = 100/2
x = 50
check:
50 + (50 -6) = 94
50 + (44) = 94
94 = 94